# Learning policies for resource allocation in business processes

### v2

@@ 5.2.1 Deep Reinforcement Learning
During training, we decrease the learning rate linearly with the total training steps to encourage exploration at the start and exploitation of the policy at the end of training. The discount factor determines the importance of future rewards. We found a high value to be optimal to account for the long-term effects of actions on the cycle time and sparse rewards. We evaluated the current model during training and saved the best model found.
In the N-network, the DRL algorithm converged to a policy where the agent only postpones its decisions. Each case in this scenario only contains a single activity before the case is completed, and the agent receives a reward. Because of this, the agent learned to prioritize one of the two activities and get high rewards by completing that activity quickly. To accomplish this, the agent learned to postpone instead of allocating a resource for the other activity. While this policy may result in high rewards, it is obviously not beneficial for the mean cycle time. To mitigate the bias induced by the reward function, we impose a penalty of -0.1 every time the agent uses the postpone action in this scenario. Penalizing the agent in this scenario led to competitive results with the best-performing benchmarks, which will be shown in Section 5.4.
We use the maskable PPO implementation of the contributory version of the popular DRL Python library Stable-Baselines3  [^1] https://github.com/Stable-Baselines-Team/stable-baselines3-contrib and create our environment using Gymnasium  [^2] https://gymnasium.farama.org/index.html. Untuned hyperparameters were kept at their default values.
@@ 5.2.2 Score-based value function approximation
To train the weights of the SVFA method, we use Bayesian optimization. We define a search space in which we optimize the value function’s weights for each business process. In our setup, we use $0\leq w_{i}\leq 100$, for $i\in\{1,\dots,7\}$. The optimization procedure runs for a maximum of 1000 trials and converges before using all trials in all business processes. The method was implemented in Python using the scikit-optimize package [^3] https://scikit-optimize.github.io/stable/.
@@ 5.3 Evaluation protocol and benchmarks
Each method is evaluated on 100 simulations of 5000 time units in each business process. The cycle time is calculated according to Equation 1. Uncompleted cases at the end of the process are truncated and included in the average cycle time calculation. This truncation ensures that the best policy is not always to postpone, as otherwise, no case completions would result in a zero average cycle time. We test if the method with the lowest cycle time significantly differs from the other methods using a T-test for two independent samples ($p\geq 0.05$).
To benchmark our methods, we compare them with three traditional heuristics and two methods from the literature park_prediction-based_2019 ; zbikowski_deep_2023 . Not all benchmarks were directly implementable for our problem, or no source code was available, therefore, we had to make small changes from the original implementation. The benchmarks and their implementation details are:
    Shortest Processing Time (SPT): we evaluate all possible assignments at a decision step on their expected processing time and choose the assignment(s) with the shortest processing time.
@@
    Random: a random assignment is made from all possible assignments at a decision step. This heuristic serves as a baseline.
@@
    Park and Song park_prediction-based_2019 : this method uses a prediction model to predict the next activity and the processing time of possible assignments. Since we do not have an event log to train a prediction model, we derive the next activity and the expected processing time of each assignment from the business process model. Therefore, the predictions we passed to the scheduling stage are 100% accurate.
@@
We chose these benchmarks because their methods are closely related to our methods and are suitable for the online scheduling problem. Considering we do not need to make any predictions, the minimum-cost maximum-flow algorithm by Park and Song park_prediction-based_2019 is an advanced version of SPT and, therefore, a relevant benchmark. We implemented RLRAM huang_reinforcement_2011 as a preliminary benchmark. However, the method does not scale well to business processes containing much stochasticity like those considered in this paper. Our implementation of RLRAM failed to converge on competitive policies within a reasonable time frame. Our implementation of RLRAM algorithm can be found in our code. Other methods from the literature, mine features from an event log to solve the resource allocation problem firouzian_cycle_2019 ; xu_resource_2008 ; zhao_optimization_2015 ; arias_framework_2016 ; kuchar_automatic_2016 . However, our evaluation framework uses simulated data without a real-world event log; therefore, these event log-based approaches could not be directly replicated.
5.4 Results
Table 2: Mean cycle time and half-width of the 95%-confidence interval for the six scenarios for DRL and SVFA, compared with three heuristics and two existing methods. Results were obtained over 100 simulations. A lower cycle time is better. The best result is in bold, and multiple bold results indicate they are not significantly different from the best method.
Business process DRL SVFA SPT FIFO Random Park & Song park_prediction-based_2019 Żbikowski et al. zbikowski_deep_2023
Low utilization 5.7 (0.09) 5.6 (0.08) $5.8\,(0.08)$ $5.9\,(0.11)$ $6.6\,(0.14)$ $6.0\,(0.09)$*
High utilization 18.9 (0.91) 18.1 (0.94) 18.7 (0.89) $27.0\,(2.77)$ $35.8\,(4.20)$ 19.4 (1.01) -*
Slow server 12.2 (0.38) 11.4 (0.31) $27.1\,(1.50)$ $19.3\,(1.33)$ $21.9\,(1.49)$ $20.5\,(1.26)$ $58.9\,(5.60)$
Slow downstream 9.9 (0.33) 9.6 (0.25) $15.4\,(0.71)$ 9.8 (0.35) $11.9\,(0.53)$ $14.7\,(0.59)$*
N-network $5.9\,(0.11)$ 5.7 (0.09) $7.4\,(0.20)$ $5.9\,(0.11)$ $6.5\,(0.16)$ $6.3\,(0.14)$ $6.7\,(0.11)$
Parallelism $11.3\,(0.38)$ $12.0\,(0.38)$ $13.8\,(0.57)$ 9.3 (0.31) $11.0\,(0.45)$ $13.7\,(0.60)$*
As can be seen in Table 2, the DDQN method by Żbikowski et al. zbikowski_deep_2023 could not converge to a competitive policy in some scenarios. The reason for this is that this algorithm can get stuck in a loop where it only selects infeasible actions. In some cases, the representation of their state space does not change until an assignment is made (see zbikowski_deep_2023 ). If the deterministic policy takes an infeasible action in that state, no assignments are made, and thus, the state never changes, resulting in indefinite postponement.
Table 2 shows that both of our methods outperform or are competitive across benchmarks, except for the scenario involving parallelism. In the low utilization scenario, strategic decisions like idling or prioritizing case completions become more viable as the load on the resources is. Our methods excel in learning these strategies. On the other hand, in the high utilization scenario, prioritizing the assignment of activity instances to the fastest available resources is crucial to prevent excessive queue buildup. This strategy aligns closely with the SPT heuristic and with Park and Song’s method park_prediction-based_2019 , as shown in these methods’ performance. Our methods can also learn this policy effectively.
In the slow server scenario, resource $r_{5}$ is inefficient in executing activity E and avoiding this assignment whenever possible is advisable (see Fig. 2(c)). Hence, strategically idling to await a faster resource’s availability may be beneficial. Park and Song’s method park_prediction-based_2019 also accommodates strategic idling, which is evident in its improved performance over SPT. Both of our methods successfully learned policies that outperform all benchmarks.
In the slow downstream scenario, the processing time of an activity is equal across resources. Therefore, the emphasis shifts from choosing what resource to allocate to what activity to execute. Finishing cases as quickly as possible leads to the lowest cycle time, as seen in the FIFO heuristic’s high performance. Our methods are also capable of learning this policy.
In the N-network, only resource $r_{10}$ can execute activity I (see Fig. 2(e)); therefore, reserving it for that activity is an effective strategy. While SVFA was able to find a policy that outperforms the other methods, our DRL model struggled. The DRL agent learned only to execute activity J and postpone otherwise, leading to high immediate rewards. However, this policy resulted in a large queue building up at activity I, and, therefore, a high average cycle time. After introducing a penalty of $-0.1$ for postponing in this scenario, the agent converged to a better policy.
In the parallelism scenario, a FIFO policy yielded the best performance. FIFO completes all activities of a case before initiating the next case, resulting in less waiting time and lower cycle times. In contrast, other methods lack information about what activity instances lead to case completions, making finding an optimal resource allocation policy more challenging. The results for the composite business processes are shown in Table 3.
Table 3: Mean cycle time and half-width of the 95%-confidence interval for composite business processes for DRL and SVFA, compared with three heuristics and two existing methods. Results were obtained over 100 simulations. A lower cycle time is better. The best result is in bold, and multiple bold results indicate they are not significantly different from the best method.
Business process DRL SVFA SPT FIFO Random
Composite 64.7 (2.63) $72.1\,(2.34)$ $98.8\,(3.80)$ $69.0\,(3.36)$ $84.9\,(4.00)$
Composite reversed 62.3 (1.77) $73.6\,(2.44)$ $106.6\,(5.18)$ $68.1\,(2.61)$ $85.0\,(4.74)$
Composite parallel 23.4 (0.92) $34.1\,(1.29)$ $35.1\,(1.73)$ $31.0\,(2.54)$ $37.1\,(2.6)$
Similarly, as seen in some of the scenarios (Table 2), the method by Żbikowski et al. zbikowski_deep_2023 converged to a policy that postponed indefinitely. Furthermore, the inference time of the method by Park and Song park_prediction-based_2019 was intractable; therefore, we could not evaluate the method for the composite business processes. This limitation was also indicated by the authors. For this reason, the results of these two benchmarks are not included in the evaluation and Table 3.
Table 3 shows that DRL outperforms all methods in the composite business processes. The performances of the composite and composite reverse business processes are similar, which means that the order of scenarios does not significantly impact the performance. This was to be expected because the scenarios in the composite business processes are still somewhat isolated, as each scenario has its own resources. While FIFO outperformed the other methods in the parallel scenario, as seen in Table 2, it could not effectively handle the dynamics of all scenarios combined, as seen in Table 3.
SVFA performed well in the scenarios but could not find competitive policies in the composite business processes. A possible explanation is that there are competing feature weights for each scenario. For example, as can be seen in Table 2, an SPT policy yields good performance in the high utilization scenario, from which we conclude that the weight of the feature $MeanAssignment$ should be high for this scenario. On the other hand, SPT makes poor assignments in the slow downstream scenario, which means the weight of $MeanAssignment$ should be low in this scenario. A balance must be found in the composite business processes for this feature to account for both scenarios, which can result in a sub-optimal policy.

### v3

@@ 5.2.1 Deep Reinforcement Learning
During training, we decrease the learning rate linearly with the total training steps to encourage exploration at the start and exploitation of the policy at the end of training. The discount factor determines the importance of future rewards. We found a high value to be optimal to account for the long-term effects of actions on the cycle time and sparse rewards. We evaluated the current model during training and saved the best model found. We train each model until the policy converges or until a maximum of 2e7 training steps (i.e., decision steps). However, we found that each policy converged much earlier than the maximum number of steps.
We use the maskable PPO implementation of the contributory version of the popular DRL Python library Stable-Baselines3  [^1] https://github.com/Stable-Baselines-Team/stable-baselines3-contrib and create our environment using Gymnasium  [^2] https://gymnasium.farama.org/index.html. Other hyperparameters were kept at their default values.
@@ 5.2.2 Score-based value function approximation
To train the weights of the SVFA method, we use Bayesian optimization. We define a search space in which we optimize the value function’s weights for each business process. In our setup, we use $0\leq w_{i}\leq 100$, for $i\in\{1,\dots,7\}$. The optimization procedure runs for a maximum of 20 trials. Each trial contains 5000 simulations of the business process to estimate the mean cycle time given the current weights. The method was implemented in Python using the scikit-optimize package [^3] https://scikit-optimize.github.io/stable/.
@@ 5.3 Evaluation protocol and benchmarks
Each method is evaluated on 100 simulations of 5000 time units in each business process. The cycle time is calculated according to Def. 14. Uncompleted cases at the end of the process are truncated and included in the average cycle time calculation. This truncation ensures that the best policy is not always to postpone, as otherwise, no case completions would result in a zero average cycle time. For each method and benchmark, we will report the mean cycle time and the 95% confidence interval (CI). A Student’s t-test for two independent samples ($p\geq 0.05$) will be performed between the best-performing method and the other methods. The best-performing method and methods not significantly different will be reported in bold.
We evaluated our methods and will present our results around the proposed arrival rate $\lambda=0.5$. However, to investigate the performance of our models under different loads, we evaluated each method using both different constant arrival rates and a dynamic arrival pattern. We trained and evaluated each model for the arrival rates $\lambda\in\{0.3,0.35,0.4,0.45,0.5,0.55,0.6\}$. In an arrival pattern, the arrival rate changes over time. Figure 4 shows the arrival pattern used derived from the BPI12W event log.
Figure 4: Arrival pattern based on the BPI12W event log
During the simulation, arrivals are sampled using the KDE curve. The pattern repeats every 250-time units, and the night time of the original event log is ignored, as our resources are continuously available. If we include this downtime in arrivals, many cases will arrive during the day and must be processed at night. The pattern is scaled such that the average arrival rate is 0.5. The pattern is implemented as follows. After each arrival, a new arrival is sampled using the maximum arrival rate in the KDE curve ($\lambda_{max}=0.88)$. This arrival is then accepted with a probability proportional to the current arrival rate. For example, if $\lambda_{current}=0.5$, then $P(accept)=\frac{\lambda_{current}}{\lambda_{max}}=0.44$.
To benchmark our methods, we compare them with three traditional heuristics and two methods from the literature [5, 8]. Some of the considered benchmarks were not directly implementable, or no source code was available. Therefore, we had to make some small changes or assumptions. The benchmarks and their implementation details are:
    Shortest Processing Time (SPT): all possible assignments at a decision step are evaluated at a decision step, and the assignment with the lowest expected processing time is chosen.
@@
    Random: a random assignment is made from all possible assignments at a decision step.
@@
    Park and Song [5]: this method uses a prediction model to predict the next activity and the processing time of possible assignments. We do not have an event log to train a prediction model, but can derive the next activity and the expected processing times from the business process model. Therefore, the predictions we passed to the scheduling stage are 100% accurate.
    Żbikowski et al. [8]: this method uses a double deep Q-network (DDQN) algorithm to train an agent to make assignments. As our business processes are larger than those evaluated in [8], we train the DDQN for more steps, equal to the number of steps we train our models. We use the same values for the number of exploration episodes and replay buffer as in the DDQN’s implementation.
We chose these benchmarks because their methods are closely related to our methods and are suitable for the online scheduling problem. Considering we do not need to make any predictions, the minimum-cost maximum-flow algorithm by Park and Song [5] is an advanced version of SPT and, therefore, a relevant benchmark. We implemented RLRAM [6]; however, the method did not converge to a competitive policy in any of our business processes. Our implementation of RLRAM algorithm can be found in our repository. Other methods from the literature mine features from an event log to solve the resource allocation problem [7, 14, 15, 16, 17]. However, our evaluation framework uses simulated data without a real-world event log; therefore, these event log-based approaches could not be replicated.
5.4 Results
In this section, we present and discuss the results of our methods compared to the benchmarks. Section 5.4.1 presents the results of the scenarios and composite business processes using a constant arrival rate of $\lambda=0.5$. Section 5.4.2 presents the results when using a range of constant arrival rates and a dynamic arrival pattern.
5.4.1 Evaluation of the business processes
This section shows the results of our methods compared with the benchmarks for the scenarios and composite business processes. Table 2 shows the results of our methods on the six scenarios. The method by Żbikowski et al. [8] can select infeasible actions, after which the state changes and another action is selected. However, in some situations, the state does not change anymore, and the policy gets stuck, selecting infeasible actions. For these scenarios, we do not report the results for this method.
Table 2: Mean cycle time and CIs for the six scenarios ($\lambda=0.5$).
Business process DRL SVFA SPT FIFO Random Park & Song [5] Żbikowski et al. [8]
Low utilization 5.8 (0.10) 5.9 (0.10) 5.9 (0.09) 6.0 (0.11) 6.5 (0.13) 6.0 (0.09) -a
High utilization 17.9 (0.72) 19.4 (0.99) 19.4 (0.96) 26.5 (1.86) 33.2 (3.07) 19.4 (1.01) -a
Slow server 11.5 (0.33) 14.7 (0.45) 26.6 (1.88) 20.8 (1.86) 21.2 (1.25) 20.5 (1.26) 58.9 (5.60)
Slow downstream 10.0 (0.35) 10.0 (0.32) 14.9 (0.61) 9.9 (0.32) 11.5 (0.39) 14.7 (0.59) -a
N-network 6.8 (0.20)* 6.1 (0.11) 7.1 (0.21) 6.0 (0.12) 6.5 (0.15) 6.3 (0.14) 6.7 (0.11)
Parallel 10.8 (0.37) 12.6 (0.54) 14.1 (0.6) 9.8 (0.35) 11.1 (0.49) 13.7 (0.60) -a
    Policy trained with -0.1 penalty for postponing
  • a
Table 2 shows that both of our methods outperform or are competitive across benchmarks, except for the scenario involving parallelism. In the low utilization scenario, strategic decisions like idling or prioritizing case completions become more viable as the resource load is lower. Our methods can effectively learn these policies. On the other hand, in the high utilization scenario, prioritizing the assignment of activity instances to the fastest available resources is crucial to prevent excessive queue buildup. This strategy aligns closely with the SPT heuristic and with Park and Song’s method [5], as shown in these methods’ performance. Our DRL method found a policy that outperforms the other methods while SVFA’s policy was competitive with the best-performing heuristics.
In the slow server scenario, resource $r_{5}$ is inefficient in executing activity E (Figure 3(c)), and this assignment should be avoided in most situations. Hence, strategically idling to await a faster resource’s availability may be beneficial. Park and Song’s method [5] also accommodates strategic idling, which is evident in its improved performance over SPT. Both of our methods successfully learned policies that outperform all benchmarks.
In the slow downstream scenario, the processing time of an activity is equal across resources. Therefore, the emphasis shifts from choosing what resource to allocate to what activity to execute. In fact, in this scenario, it is optimal to always perform the activity H if possible, as it completes the case and stops the cycle time. While this assignment is not more efficient on the short-term, it must be completed and thus should be prioritized for case completion. This policy is identical to FIFO’s policy, which explains its good performance. Our methods are also capable of learning this policy.
In the N-network, only resource $r_{10}$ is eligible to execute activity I (see Figure 3(e)); therefore, reserving it for that activity is an effective strategy. While SVFA was able to find a policy that outperforms the other methods, our DRL model struggled. The DRL agent learned only to execute activity J and postpone otherwise, leading to high immediate rewards. However, this policy resulted in a large queue building up at activity I, resulting in a high average cycle time. After introducing a penalty of $-0.1$ for postponing in this scenario, the agent converged to a better policy.
Similar to scenario 4 (Figure 3(d)), the parallelism scenario contains two activities with identical processing times across the two different activities. As both activities must be executed before a case is completed, it is again optimal to assign available resources to the longest waiting case, which is what the FIFO policy does. In contrast, our methods and other benchmarks do not have explicit information about the previously executed activities of a specific case, meaning that no distinction can be made between new cases and partially completed cases. This lack of information makes finding the optimal policy challenging for this scenario, which is reflected by the results.
While these scenarios may not capture all possible process patterns or flows that can occur in a business process, they capture the most relevant dynamics in terms of possible control-flow relations, resource loads, and authorization structures. For example, the utilization-based scenarios can represent different resource loads that can change as the process executes. Furthermore, the slow server and N-network emulate different resource authorization structures, and the N-network and parallel scenario model the parallel (AND) and exclusive (XOR) control-flow patterns. While we test our models on these scenarios, there are other scenarios, such as a W-network, which can also be investigated. To illustrate the robustness of our conclusions in other scenarios, we will now extend our evaluation to composite business processes. In addition, in another paper, we applied our method to real-world business processes, containing dynamics such as re-entrant flows and the temporal availability of resources [33].
In the second part of the evaluation, we evaluated the three composite business processes. Figure 5 shows the progression of total reward and mean cycle time per episode during training with our DRL models. The policies for the two sequential processes converge quickly, stabilizing after approximately 100 episodes. The policy for the composite parallel process takes around 200 episodes to converge due to its increased complexity. In this process, various activities can lead to case completion, unlike sequential processes where only the final activity completes the case. All three composite processes converged before reaching the allowed training time (maximum of 2e7 decision steps is approximately 1800 episodes for these processes). In our SVFA method, each model used the 20 allowed trials.
Figure 5: Total reward and mean cycle time during training of DRL models.
The results for the composite business processes using a constant arrival rate of $\lambda=0.5$ are shown in Table 3.
Table 3: Mean cycle time and CIs for the composite processes ($\lambda=0.5$).
Business process DRL SVFA SPT FIFO Random
Composite 64.7 (2.63) 72.9 (2.36) 100.9 (4.07) 69.7 (3.5) 86.5 (4.12)
Composite reversed 61.5 (1.50) 73.8 (2.48) 110.7 (4.77) 70.0 (3.7) 88.0 (4.53)
Composite parallel 23.8 (0.99) 33.4 (1.52) 35.2 (1.71) 29.3 (1.73) 41.9 (3.99)
Similarly, as seen in some of the scenarios (Table 2), the method by Żbikowski et al. [8] converged to a policy that postponed indefinitely. Furthermore, the inference time of the method by Park and Song [5] was intractable; therefore, we could not evaluate the method for the composite business processes. This limitation was also indicated by the authors. For this reason, the results of these two benchmarks are not included in the evaluation and Table 3.
Table 3 shows that DRL consistently outperforms all other methods in composite business processes. The performance of both composite and composite reverse business processes is comparable, indicating that the sequence of scenarios does not significantly affect the outcomes. This result is expected, as the scenarios within the composite business processes remain relatively isolated, each utilizing its own resources. Although FIFO excelled in the parallel scenario, as shown in Table 2, it struggled to handle the complexities of combined scenarios, as shown in Table 3.
While SVFA performed well in the scenarios, its relative performance decreased in the composite processes. In the two sequential composite processes, SVFA showed similar performance, but in the parallel composite process, SVFA performed slightly worse than FIFO. However, SVFA still outperformed SPT, showcasing that this method is adaptable to larger process instances. A possible explanation is that there are competing feature weights for each scenario. For example, as can be seen in Table 2, an SPT policy yields good performance in the low utilization scenario, which means that the weight of the feature $MeanAssignment$ should be high for this scenario. On the other hand, SPT makes poor assignments in the slow downstream scenario, which means the weight of $MeanAssignment$ should be low in this scenario. In the composite processes, this feature must be balanced to account for both scenarios, which can result in a sub-optimal policy.
5.4.2 Evaluation using varying and dynamic arrival rates
To further investigate the effectiveness of our methods, we trained and tested them using varying and dynamic arrival rates. First, as discussed in Section 5.3, we trained and evaluated our methods using different constant arrival rates ranging from 0.3 to 0.6. Second, we use an arrival pattern that dynamically changes the arrival rate as the process is executed.
Figure 6 compares their performance to heuristics in the high utilization and slow server scenario with arrival rates from 0.3 to 0.6. The results show that all policies perform similarly when the process load is low. In these cases, resources are often idle, and simply assigning a resource is already a decent policy. On the other hand, as the load in the process increases, efficiently allocating resources becomes more important. Figure 6 shows that as the arrival rate increases, the policies diverge. The figure shows that the heuristics are effective for processes with a low load but become unsuitable as the process load and complexity of the assignment problem increase. Our methods, however, are still capable of learning relatively effective policies.
(a) High utilization
(b) Slow server
Figure 6: Comparison of policies in the high utilization and slow server scenarios using a range of arrival rates.
Figure 6 shows that the cycle time increases rapidly when the arrival rates exceed $\lambda=0.5$, which indicates that the system becomes unstable beyond this point. This pattern is also observable in the utilization rate of the resources. For instance, Figure 7 demonstrates the resource utilization in the slow server scenario.
Figure 7: Utilization rate of the resources in the slow server scenario.
As can be seen in Figure 7, the resource utilization approaches 1 at $\lambda=0.55$, which shows that the process indeed becomes unstable at this arrival rate as the resources are always occupied. We found that using an arrival rate of $\lambda=0.5$ provides a challenging problem while keeping the system stable. Interestingly, Figure 7 and Figure 6(b) show that a lower cycle time can be achieved with lower resource utilization, demonstrating the importance of an effective resource allocation policy in organizations. The behavior of the resources in other scenarios is the same as the one presented, where the utilization increases with the arrival rate in a similar trend.
@@ 5.4.2 Evaluation using varying and dynamic arrival rates
Table 4: Mean cycle time and CIs for the six scenarios using an arrival pattern.
[…]
Low utilization 17.5 (0.37) 17.2 (0.40) 17.9 (0.44) 16.8 (0.42) 17.4 (0.4) 21.4 (0.43)
High utilization 45.0 (1.58) 47.9 (3.07) 44.5 (1.02) 44.4 (1.33) 50.6 (2.11) 63.9 (3.73)
Slow server 35.2 (0.88) 34.6 (0.79) 39.2 (0.87) 69.9 (2.87) 41.6 (1.35) 49.8 (2.26)
Slow downstream 28.8 (0.57) 29.1 (0.64) 37.5 (0.86) 49.4 (1.15) 28.4 (0.58) 33.6 (0.64)
N-network 18.4 (0.45) 18.4 (0.38) 13.2 (0.29) 18.6 (0.53) 12.9 (0.29) 15.7 (0.37)
Parallel 35.4 (0.79) 29.4 (0.58) 43.5 (1.01) 48.2 (1.14) 28.5 (0.67) 32.7 (0.68)
[…]
The SVFA method performed slightly worse under the arrival pattern, as it is only the (tied) best-performing method in two out of the six scenarios. However, in the other scenarios, it can still learn a policy that outperforms some of the heuristics. The results indicate that for both our methods, it is more challenging to learn a good policy when using dynamic arrivals compared to when using a constant arrival rate.
In the second part of the evaluation, we evaluated our methods on the composite processes using varying and dynamic arrival rates. Figure 8 shows the performance of our methods using varying arrival rates in the composite processes. Arrival rates above $\lambda=0.5$ are excluded to improve the figure’s clarity and because the processes are unstable above this rate.
(a) Composite
(b) Composite reversed
(c) Composite parallel
Figure 8: Comparison of policies in the composite business processes using a range of arrival rates.
The figure shows that the relative performance of each method remains consistent across different arrival rates. Our DRL method consistently outperforms or matches the performance of other methods across all rates. At lower arrival rates, DRL performs similarly to the FIFO policy. This is because, under low load conditions (i.e., low arrival rates), there is minimal queuing, making policies that prioritize case completions effective regardless of the resources executing the activity instances. Both FIFO and our DRL method adopt this policy. However, as the load increases ($\lambda=0.5$), the performance of FIFO and DRL diverges. The DRL method learns to balance prioritizing case completions with efficient resource utilization, a balance that the FIFO policy fails to achieve.
In both the composite and composite reversed processes, our SVFA method performs comparably to a random policy at arrival rates below $\lambda=0.5$, and slightly worse than the best-performing methods. The random policy’s performance at lower arrival rates suggests that strategic resource allocation policies offer limited benefits in such scenarios, making a random policy potentially sufficient. However, at $\lambda=0.5$, the performance of SVFA diverges from the random policy, indicating that correct resource allocation can significantly impact cycle time.
[…]
Table 5: Mean cycle time and CIs for the composite processes using an arrival pattern.
[…]
Composite 120.9 (3.88) 98.3 (3.37) 111.1 (2.99) 128.9 (4.0) 96.8 (2.95) 111.8 (3.39)
Composite reversed 126.4 (3.76) 107.6 (3.66) 121.1 (3.75) 179.1 (5.73) 96.3 (3.51) 113.0 (3.69)
Composite parallel 63.2 (1.73) 51.5 (1.63) 85.4 (1.83) 73.1 (1.85) 55.1 (2.28) 68.6 (3.19)
Our methods, under a constant arrival rate, can learn policies that outperform or match the best benchmarks in archetypal business processes (Table 2). In all composite processes, our DRL method finds a policy with the lowest cycle time (Table 3).
[…]